Graph Linear Equations:
(2). Is (4,1) a solution to y=3x-11
(7). Is (-2,5) a solution to 3x+2y=9

Answer by jim_thompson5910(35256)

Graph Linear Equations:
(2). Is (4,1) a solution to y=3x-11

(4,1) means x = 4 and y = 1

y=3x-11
1=3*4-11 ... replace x with 4, replace y with 1
1=12-11
1=1 ... true

Since 1=1 is true, this means y=3x-11 is true when x=4,y=1. So (4,1) is indeed a solution to y=3x-11. Visually, the point (4,1) lies on the line y=3x-11

-------------------------------------------------------------------

(7). Is (-2,5) a solution to 3x+2y=9

(-2,5) means x = -2 and y = 5

3x+2y=9
3(-2)+2(5)=9 ... plug in x = -2 and y = 5
-6+10=9
4=9 ... false

The last equation is false, so the first equation is false when x = -2 and y = 5.
(-2,5) is NOT a solution to 3x+2y=9
The point (-2,5) does NOT lie on the line 3x+2y=9